I have a graf that the domaint is 0,1,2,3 and therange is 0,.5,2,4.5 what is the funtion if (1.5,y) is part of the graf
Answer by stanbon(75887)

I have a graf that the domaint is 0,1,2,3 and therange is 0,.5,2,4.5 what is the funtion if (1.5,y) is part of the graf
-----------------
Plot the three points: (0,0,5),(1,2),(2,2),(3,4.5)
It looks like a parabola:
y=ax^2+bx+c
Substitute to find a, b, and c,
If x=0 ; 0.5=0+0+c; so c=1/2
The y = ax^2+bx+1/2
If x=1; 2 = a+b+1/2; so a+b = (3/2)
If x=2; 9/2 = 4a+2b+(1/2):
1st: a+b = (3/2)
2nd: 2a+b = 2
----------
Subtract 1st from 2nd to get:
a= 1/2
----------------
Substitute into a+b=(3/2) to get
b = 1
-------------
EQUATION:
y = (1/2)x^2 + x + (1/2)
So, if x=1.5, y = (1/2)(9/4)+(3/2)+(1/2) = 9/8 + 2 = 25/8 = 3 1/8
===================
Cheers,
Stan H.